Explain when a system of equations will have no solutions and when it will have an infinite number of solutions.
Answer by Mathtut(3670)

Equations describing the same line
· have an Infinite number of solutions such as.....x+y=3 and 3x+3y=9...these are the same lines....the 2nd is 3 time the first.
:
Lines that do not intersect (Parallel Lines; having the same slope)
have No solutions
example: y=2/3x+7
........y=2/3x+9